# Underdetermined Fourier Extensions for Surface Partial Differential Equations

### v2

@@ 3. Proposed Methods & Analysis
Among existing approaches are interpolation method based on using a standard RBF basis, such as the Piret’s method [piret12]. Like their flat domain counterparts such as Kansa’s original method [kansa90], these methods are often quite successful in practice, but face some theoretical issues due to the possible singularity of the interpolation matrix once derivatives are involved. As with most RBF methods, the accuracy of these methods is often severely limited by conditioning, leading to errors much higher than machine error.
Other existing approaches are RBF-FD methods, which are particularly useful for large, time-dependent problems due to the sparsity of the differentiation matrices produced. These methods can either work combined with the Closest Point Method [petra18] or by more directly approximating the surface operators using estimates of a local level set [alvar21]. Being local methods, such methods typically have limited orders of convergence. Furthermore, stability (for time-stepping) and non-singularity (for elliptic problems) of such methods is currently unknown.
Lastly, various implementations using standard RBF basis functions and least-squares collocation have also been used. Significant analysis of these sorts of approaches, along with numerical testing, has been done by Chen and Ling [chen20]. A quasi-uniform point cloud is required, and a possibly quite dense point cloud of collocation points is required relative to density of the point cloud of RBF centres for guaranteed convergence. Alternatively, one may approximate a weak form of the PDE to achieve symmetry of the discretized Laplace-Beltrami operator, as in the work by Yan, Jiang, and Harlim [yan23], which overcomes a weakness of standard interpolation approaches, but requires knowledge of the distribution from which collocation points are sampled, which may not be available.
@@
Existing Fourier extension methods typically use an overdetermined system. Such methods for function approximation and bulk PDE problems in 2D have been recently proposed [stein17, matth18, bruno22], though convergence analysis is typically limited to 1D. Analysis of overdetermined methods also relies on relating the discrete least squares problem to an $L^{2}$-norm minimization problem through quadrature. Such an approach would not be feasible on all but a few surfaces where high-order quadrature schemes can be developed.
@@ 3.1. Hermite-Birkhoff Fourier Extension Problem
We first set up a constrained optimization problem. Let $S\subset\mathbb{R}^{m}$ be our domain of interest and let $\Omega\supset S$ be a box. Select a point cloud $S_{\tilde{N}}:=\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}\subset S$ and linear differential operators $\left\{\mathcal{F}_{k}\right\}_{k=1}^{\tilde{N}}$ of maximum order $p\in\mathbb{N}$, where $\mathcal{F}_{k}=\sum_{\left|\alpha\right|\leq p}c_{k,\alpha}\partial^{\alpha}$ and the coefficient functions $c_{k,\alpha}$ are bounded and defined in a neighbourhood of $\boldsymbol{x}_{k}$. Let $\left\{\phi_{n}\right\}_{n=1}^{\infty}$ be the Fourier basis for the box consisting of complex exponential eigenfunctions of the Lapacian on $\Omega$ so that $\phi_{n}\left(\boldsymbol{x}\right)=e^{i\boldsymbol{\omega}_{n}\cdot\boldsymbol{x}}$ for some frequency $\boldsymbol{\omega}_{n}$. Let $d>0$ be a sequence and let $\left\{f_{k}\right\}_{k=1}^{\tilde{N}}\subset\mathbb{C}$ be interpolation values. We then consider the problem:
@@
If the feasible set for (3.1) is non-empty and $d^{-\frac{1}{2}}\left\|\boldsymbol{\omega}\right\|_{2}^{p}\in\ell^{2}$, where we use the notation $\left\|\boldsymbol{\omega}\right\|_{2}^{p}$ to mean the sequence $\left\{\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p}\right\}_{n=1}^{\infty}$, then the solution to (3.1) is unique.
@@
If $d^{-\frac{1}{2}}\left\|\boldsymbol{\omega}\right\|_{2}^{p}\in\ell^{2}$, then noting that each $\mathcal{F}_{k}$ is of order at most $p$ and has bounded coefficient functions, it is clear that $\mathcal{L}_{\text{HB}}$ is bounded, since for $q\in\left\{0,1,\ldots,p\right\}$,
[…]
Since $\mathcal{L}_{\text{HB}}$ is bounded, $\mathcal{N}\left(\mathcal{L}_{\text{HB}}\right)$ is closed and the constraint set is closed and convex. By a standard result in functional analysis (see, for example, 3.3-1 of Kreyszig [kreys91]), if the constraint set is non-empty, there is a unique element with a minimum norm; this is the unique solution. ∎
@@ 3.2. Native Hilbert Spaces
For example, if $d=\left(1+\left\|\boldsymbol{\omega}\right\|_{2}^{2}\right)^{q}$, then $\mathcal{H}\left(d\right)=H^{q}\left(\Omega\right)$; this would be a form of minimum Sobolev norm interpolation [chand13, chand18], which has been used for PDEs in flat domains previously in a different setup [chand15]. Note that we often refer to $\left\|\cdot\right\|_{d}$ as the $d$-norm.
@@ 3.3. The Adjoint Range and Connection to Radial Basis Functions
Thus, $\boldsymbol{\Phi}$ is positive semi-definite. Moreover, $\boldsymbol{\alpha}^{*}\boldsymbol{\Phi}\boldsymbol{\alpha}\neq 0$ due to the existence of an interpolant (namely, the norm-minimizing interpolant in $\mathcal{H}\left(d\right)$) in $\text{span}\left\{\psi_{k}\right\}_{k=1}^{\tilde{N}}$ for any set of interpolation values $\left\{f_{i}\right\}_{i=1}^{\tilde{N}}$. This implies uniqueness of $\boldsymbol{\beta}$ since the system is square. That is, for any set of interpolation conditions $\left(\mathcal{F}_{i}\tilde{u}\right)\left(\boldsymbol{x}_{i}\right)=f_{i}$ for $i\in\left\{1,2,\ldots,\tilde{N}\right\}$, there is a unique, $d$-norm minimizing interpolant in $\text{span}\left\{\psi_{k}\right\}_{k=1}^{\tilde{N}}$ such that the system to find the coefficients of the interpolant in the $\left\{\psi_{k}\right\}_{k=1}^{\tilde{N}}$ basis is self-adjoint and positive definite.
Also note that $\left(\psi_{k},\psi_{j}\right)_{d}=\left(\mathcal{L}_{\text{HB}}^{*}\boldsymbol{e}_{k},\mathcal{L}_{\text{HB}}^{*}\boldsymbol{e}_{j}\right)_{\ell^{2}}=\boldsymbol{\Phi}_{kj}$. This is closely related to symmetric Hermite RBF methods for PDEs [sun94, frank98], which have been applied to problems in flat domains and rely on a similar norm-minimizing property.
@@ 3.4. Setting up a PDE
We now set up a PDE problem. Let $\Omega^{\left(j\right)}\subseteq\Omega$ for $j\in\left\{1,2,\ldots,N_{\Omega}\right\}$ all have co-dimension zero and let $S^{\left(j\right)}\subset\Omega$ for $j\in\left\{1,2,\ldots,N_{S}\right\}$ all be manifolds of co-dimension greater than zero or single points, where $\Omega$ is a box-shaped domain in $\mathbb{R}^{m}$. Define point clouds $\Omega_{\tilde{N}_{j}}^{\left(j\right)}:=\left\{\boldsymbol{x}_{k}^{\left(j\right)}\right\}_{k=1}^{\tilde{N}_{j}}\subset\Omega^{\left(j\right)}$ and $S_{\tilde{M}_{j}}^{\left(j\right)}:=\left\{\boldsymbol{y}_{k}^{\left(j\right)}\right\}_{k=1}^{\tilde{M}_{j}}\subseteq S^{\left(j\right)}$. We consider the problem:
(3.4) $\mathcal{F}^{\left(j\right)}u\biggr{|}_{\Omega^{\left(j\right)}}$ $=f^{\left(j\right)}\text{, for each $j\in\left\{1,2,\ldots,N_{\Omega}\right\}$}$
$\mathcal{G}^{\left(j\right)}u\biggr{|}_{S^{\left(j\right)}}$ $=g^{\left(j\right)}\biggr{|}_{S^{\left(j\right)}}\text{, for each$j\in\left\{1,2,\ldots,N_{S}\right\}$},$
where $\mathcal{F}^{\left(j\right)}$ and $\mathcal{G}^{\left(j\right)}$ are linear differential operators of order at most $p$, $f^{\left(j\right)}$ is a function defined on $\Omega^{\left(j\right)}$, and $g^{\left(j\right)}$ is defined in a neighbourhood of $S^{\left(j\right)}$ for each $j$. Note that we can extend functions on $S^{\left(j\right)}$ to an open neighbourhood of $S^{\left(j\right)}$ while preserving their differentiability using Proposition 5, as long as $S^{\left(j\right)}$ is $C^{2}$; we note this here since the degree of differentiability of $g^{\left(j\right)}$ shows up in convergence rates in Subsection 3.5. This form covers all strong-form linear PDEs on piecewise smooth manifolds we may consider. We then have the discretization:
@@
subject to: $\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}b_{n}\left(\mathcal{F}^{\left(j\right)}\phi_{n}\right)\left(\boldsymbol{x}_{k}^{\left(j\right)}\right)=f^{\left(j\right)}\left(\boldsymbol{x}_{k}^{\left(j\right)}\right)\text{, foreach $j,k$}$
$\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}b_{n}\left(\mathcal{G}^{\left(j\right)}\phi_{n}\right)\left(\boldsymbol{y}_{k}^{\left(j\right)}\right)=g^{\left(j\right)}\left(\boldsymbol{y}_{k}^{\left(j\right)}\right)\text{, foreach $j,k$}.$
@@ 3.5. Analysis for Hermite-Birkhoff Problem
Let $h_{\text{max}}$ be the largest fill distance of each point cloud $\Omega_{\tilde{N}_{j}}^{\left(j\right)}$ and $S_{\tilde{M}_{j}}^{\left(j\right)}$ in their respective domains $\Omega^{\left(j\right)}$ and $S^{\left(j\right)}$. We can then prove uniform convergence of $\mathcal{F}^{\left(j\right)}\tilde{u}\to f^{\left(j\right)}$ and $\mathcal{G}^{\left(j\right)}\tilde{u}\to g^{\left(j\right)}$; convergence is high-order or super-algebraic with respect to $h_{\text{max}}$ as long as certain smoothness requirements are met. This, or similar arguments with different norms, is enough to prove high-order convergence of $\tilde{u}$ to a solution to the PDE if the PDE is well-posed with respect to the norms in consideration. Proposition 12 will later show convergence to a solution of the PDE directly under weaker conditions, albeit without rate estimates.
Note for the following results that we typically omit the notation for the restriction of a function when the restricted domain is otherwise clear, such as in a norm. For example, we may write $\left\|\tilde{u}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}$ to mean $\left\|\tilde{u}|_{S^{\left(j\right)}}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}$. We also note again here that we use $\left\|\boldsymbol{\omega}\right\|_{2}$ to mean the sequence $\left\{\left\|\boldsymbol{\omega}_{n}\right\|_{2}\right\}_{n=1}^{\infty}$ (it is a sequence of finite dimensional 2-norms, not an $\ell^{2}$ norm).
@@
Let $\tilde{b}$ be the solution to (3.5) and let $\tilde{u}=\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{n}\phi_{n}$. Assume $u\in\mathcal{H}\left(d\right)$ is any solution to (3.4). Let $j\in\left\{1,2,\ldots,N_{\Omega}\right\}$ be fixed. Assume $\Omega^{\left(j\right)}$ satisfies an interior cone condition, and assume $q\in\mathbb{N}$ is such that $f^{\left(j\right)}\in H^{q}\left(S\right)$ and $F^{\left(j,q\right)}d^{-\frac{1}{2}}\in\ell^{2}$, where $F^{\left(j,q\right)}:=\left\{\left\|\mathcal{F}^{\left(j\right)}\phi_{n}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}\right\}_{n=1}^{\infty}$. Then, for small enough $h_{\text{max}}$, there exists a constant $B_{q}>0$ such that
$\left\|\mathcal{F}^{\left(j\right)}\tilde{u}-f^{\left(j\right)}\right\|_{L^{\infty}\left(\Omega^{\left(j\right)}\right)}\leq B_{q}h_{\text{max}}^{q-\frac{1}{2}-\frac{m}{2}}\left(C_{j,q}\left\|u\right\|_{d}+\left\|f^{\left(j\right)}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}\right),$
where $C_{j,q}:=\left\|F^{\left(j,q\right)}d^{-\frac{1}{2}}\right\|_{\ell^{2}}$.
@@
We have:
$\left\|\mathcal{F}^{\left(j\right)}\tilde{u}-f^{\left(j\right)}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}$ $\leq\left\|\mathcal{F}^{\left(j\right)}\tilde{u}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}+\left\|f^{\left(j\right)}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}$
$\leq\sum_{n=1}^{\infty}\left\|\mathcal{F}^{\left(j\right)}\phi_{n}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}d_{n}^{-\frac{1}{2}}\left|\tilde{b}_{n}\right|+\left\|f^{\left(j\right)}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}$
[…]
where we note $u$ is a feasible Hermite-Birkhoff interpolant so $\left\|\tilde{b}\right\|_{\ell^{2}}=\left\|\tilde{u}\right\|_{d}\leq\left\|u\right\|_{d}$. By Corollary 7, we then have that there exists some constant $B_{q}$ such that for small enough $h_{\text{max}}$,
$\left\|\mathcal{F}^{\left(j\right)}\tilde{u}-f^{\left(j\right)}\right\|_{L^{\infty}\left(\Omega^{\left(j\right)}\right)}\leq B_{q}h_{\text{max}}^{q-\frac{m}{2}}\left(C_{j,q}\left\|u\right\|_{d}+\left\|f^{\left(j\right)}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}\right)\text{, as required}.$
We note that the $F^{\left(j,q\right)}d^{-\frac{1}{2}}\in\ell^{2}$ condition tells us how quickly our choice of $d^{-\frac{1}{2}}$ must decay. If $\mathcal{F}^{\left(j\right)}$ is order $p$ and has bounded coefficient functions with $q$ bounded derivatives, then $F^{\left(j,q\right)}$ will grow as $\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p+q}$ as $n\to\infty$.
The error on manifolds of co-dimension greater than zero is slightly more involved since the boundary does not satisfy an interior cone condition. Noting that if $S^{\left(j\right)}$ is a point, then $\mathcal{G}^{\left(j\right)}\tilde{u}=g^{\left(j\right)}$ on that point as long as $S^{\left(j\right)}_{\tilde{M}_{j}}$ is non-empty, we now cover the case that $S^{\left(j\right)}$ has dimension at least one.
Proposition 10.
Assume $S^{\left(j\right)}$ is of dimension $m_{j}>1$ and can be parametrized by a finite atlas $\left\{\sigma^{\left(j\right)}_{k}\right\}_{k=1}^{M}$ where $\sigma^{\left(j\right)}_{k}:U_{k}^{\left(j\right)}\to S^{\left(j\right)}_{k}$ and each $U_{k}^{\left(j\right)}\subset\mathbb{R}^{m_{j}}$ satisfies an interior cone condition. Let $q\in\mathbb{N}$ and assume each $\sigma_{k}^{\left(j\right)}$ is $C^{q}$ with bounded derivatives up to order $q$. Also assume that each $\sigma_{k}^{\left(j\right)}$ has an inverse metric tensor with a bounded norm on $\sigma_{k}^{\left(j\right)}\left(U_{k}^{\left(j\right)}\right)$. Let $j\in\left\{1,2,\ldots,N_{S}\right\}$ be fixed. Also assume $g^{\left(j\right)}$ has bounded derivatives up to order $q$ and $G^{(j,q)}d^{-\frac{1}{2}}\in\ell^{2}$, where $G^{(j,q)}:=\left\{\max_{\left|\beta\right|\leq q}\left\|\partial^{\beta}\mathcal{G}^{\left(j\right)}\phi_{n}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right\}_{n=1}^{\infty}$. Then, for small enough $h_{\text{max}}$, there exists a constant $\tilde{B}_{q}>0$ such that
$\left\|\mathcal{G}^{\left(j\right)}\tilde{u}-g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\leq\tilde{B}_{q}h_{\text{max}}^{q-\frac{m_{j}}{2}}\left(\tilde{C}_{j,q}\left\|u\right\|_{d}+\max_{\left|\beta\right|\leq q}\left(\left\|\partial^{\beta}g\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right)\right),$
where $u\in\mathcal{H}\left(d\right)$ is again a solution to (3.4) and $\tilde{C}_{j,q}:=\left\|G^{\left(j,q\right)}d^{-\frac{1}{2}}\right\|_{\ell^{2}}$.
Proof.
@@
Since the derivatives of $\sigma^{\left(j\right)}_{k}$ are bounded, repeated application of chain rule shows there exist constants $A_{k,q}$ so that for each multi-index $\alpha$ of order less than or equal to $q$,
@@
$\ \left\|\partial^{\alpha}\left(\mathcal{G}^{\left(j\right)}\tilde{u}\circ\sigma^{\left(j\right)}_{k}-g^{\left(j\right)}\circ\sigma^{\left(j\right)}_{k}\right)\right\|_{L^{\infty}\left(U^{(j)}_{k}\right)}$
$\leq$ $\ A_{k,q}\max_{\left|\beta\right|\leq q}\left\|\partial^{\beta}\left(\mathcal{G}^{\left(j\right)}\tilde{u}-g^{\left(j\right)}\right)\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}$

### v3

@@ 3. Proposed Methods & Analysis
Among existing approaches are interpolation methods based on using a standard RBF basis, such as the method of Piret (2012). Like their flat domain counterparts such as Kansa’s original method (see Kansa, 1990), these methods are often quite successful in practice but face some theoretical issues due to the possible singularity of the interpolation matrix once derivatives are involved. As with most RBF methods, the accuracy of these methods is often severely limited by conditioning, leading to errors much higher than machine error.
Other existing approaches are RBF-FD methods, which are particularly useful for large, time-dependent problems due to the sparsity of the differentiation matrices produced. These methods can either work combined with the Closest Point Method (Petras et al., 2018) or by more directly approximating the surface operators using estimates of a local level set (Álvarez et al., 2021). Being local methods, such methods typically have limited orders of convergence. Furthermore, stability (for time-stepping) and non-singularity (for elliptic problems) of such methods are currently unknown.
Lastly, various implementations using standard RBF basis functions and least-squares collocation have also been used. Significant analysis of these sorts of approaches, along with numerical testing, has been done by Chen and Ling (2020). A quasi-uniform point cloud is required, and a possibly quite dense point cloud of collocation points is required relative to density of the point cloud of RBF centres for guaranteed convergence. Alternatively, one may approximate a weak form of the PDE to achieve symmetry of the discretized Laplace-Beltrami operator, as in the work by Yan et al. (2023), which overcomes a weakness of standard interpolation approaches, but requires knowledge of the distribution from which collocation points are sampled, which may not be available.
@@
Existing Fourier extension methods typically use an overdetermined system. Such methods for function approximation and bulk PDE problems in 2D have been recently proposed (see, for example, Stein et al., 2017; Matthysen and Huybrechs, 2018; Bruno and Paul, 2022), though convergence analysis is typically limited to 1D. Analysis of overdetermined methods also relies on relating the discrete least squares problem to an $L^{2}$-norm minimization problem through quadrature. Such an approach would not be feasible on all but a few surfaces where high-order quadrature schemes can be developed.
@@ 3.1. Hermite-Birkhoff Fourier Extension Problem
We first set up a constrained optimization problem. Let $S\subset\mathbb{R}^{m}$ be our domain of interest and let $\Omega\supset S$ be a box. Select a point cloud $S_{\tilde{N}}:=\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}\subset S$ and linear differential operators $\left\{\mathcal{F}_{k}\right\}_{k=1}^{\tilde{N}}$ of maximum order $p\in\mathbb{N}$, where $\mathcal{F}_{k}=\sum_{\left|\alpha\right|\leq p}c_{k,\alpha}\partial^{\alpha}$ and the coefficient functions $c_{k,\alpha}$ are bounded and defined in a neighbourhood of $\boldsymbol{x}_{k}$ and $\partial^{\alpha}$ are the usual partial derivatives in $\mathbb{R}^{m}$. We note here that surface differential operators can be written in this form using extensions to the box $\Omega$ (see Lemma 2 and Corollary 3). Let $\left\{\phi_{n}\right\}_{n=1}^{\infty}$ be the Fourier basis for the box consisting of complex exponential eigenfunctions of the Lapacian on $\Omega$ so that $\phi_{n}\left(\boldsymbol{x}\right)=e^{i\boldsymbol{\omega}_{n}\cdot\boldsymbol{x}}$ for some frequency $\boldsymbol{\omega}_{n}$. Let $d=\left\{d_{n}\right\}_{n=1}^{\infty}$ be a sequence where each $d_{n}>0$, and let $\left\{f_{k}\right\}_{k=1}^{\tilde{N}}\subset\mathbb{C}$ be interpolation values. We then consider the problem:
@@
If the feasible set for (3.1) is non-empty and $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}$
$\left\{\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p}d_{n}^{-\frac{1}{2}}\right\}_{n=1}^{\infty}\in\ell^{2}$, then the solution to (3.1) is unique.
@@
If $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$, then noting that each $\mathcal{F}_{k}$ is of order at most $p$ and has bounded coefficient functions, it is clear that $\mathcal{L}_{\text{HB}}$ is bounded, since for $q\in\left\{0,1,\ldots,p\right\}$,
$\left|\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}b_{n}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{q}\phi_{n}\left(\boldsymbol{x}_{k}\right)\right|\leq\left\|\left\|\boldsymbol{\omega}\right\|_{2}^{q}d^{-\frac{1}{2}}\right\|_{\ell^{2}}\left\|b\right\|_{\ell^{2}}\text{, by Cauchy-Schwarz}.$
Since $\mathcal{L}_{\text{HB}}$ is bounded, $\mathcal{N}\left(\mathcal{L}_{\text{HB}}\right)$ is closed and the constraint set is closed and convex. By a standard result in functional analysis (see, for example, Kreyszig, 1991, 3.3-1), if the constraint set is non-empty, there is a unique element with a minimum norm; this is the unique solution. ∎
@@ 3.2. Native Hilbert Spaces
For example, if $d_{n}=\left(1+\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2}\right)^{q}$, then $\mathcal{H}\left(d\right)=H^{q}\left(\Omega\right)$; this would be a form of minimum Sobolev norm interpolation (Chandrasekaran et al., 2013, 2018), which has been used for PDEs in flat domains previously in a different setup (Chandrasekaran and Mhaskar, 2015). Note that we often refer to $\left\|\cdot\right\|_{d}$ as the $d$-norm.
@@ 3.3. The Adjoint Range and Connection to Radial Basis Functions
Thus, $\boldsymbol{\Phi}$ is positive semi-definite. Moreover, $\boldsymbol{\alpha}^{*}\boldsymbol{\Phi}\boldsymbol{\alpha}\neq 0$ for $\boldsymbol{\alpha\neq 0}$ due to the existence of an interpolant (namely, the norm-minimizing interpolant in $\mathcal{H}\left(d\right)$) in $\text{span}\left\{\psi_{k}\right\}_{k=1}^{\tilde{N}}$ for any set of interpolation values $\left\{f_{i}\right\}_{i=1}^{\tilde{N}}$. This implies uniqueness of $\boldsymbol{\beta}$ since the system is square. That is, for any set of interpolation conditions $\left(\mathcal{F}_{i}\tilde{u}\right)\left(\boldsymbol{x}_{i}\right)=f_{i}$ for $i\in\left\{1,2,\ldots,\tilde{N}\right\}$, there is a unique, $d$-norm minimizing interpolant in $\text{span}\left\{\psi_{k}\right\}_{k=1}^{\tilde{N}}$ such that the system to find the coefficients of the interpolant in the $\left\{\psi_{k}\right\}_{k=1}^{\tilde{N}}$ basis is self-adjoint and positive definite.
Also note that $\left(\psi_{k},\psi_{j}\right)_{d}=\left(\mathcal{L}_{\text{HB}}^{*}\boldsymbol{e}_{k},\mathcal{L}_{\text{HB}}^{*}\boldsymbol{e}_{j}\right)_{\ell^{2}}=\boldsymbol{\Phi}_{kj}$. This is closely related to symmetric Hermite RBF methods for PDEs (see Sun, 1994; Franke and Schaback, 1998), which have been applied to problems in flat domains and rely on a similar norm-minimizing property.
@@ 3.4. Setting up a PDE
We now set up a PDE problem. Let $S^{\left(j\right)}\subset\Omega$ for $j\in\left\{1,2,\ldots,N_{S}\right\}$ all be manifolds, open bulk domains, or single points, where $\Omega$ is a box-shaped domain in $\mathbb{R}^{m}$. Define point clouds $S_{\tilde{N}_{j}}^{\left(j\right)}:=\left\{\boldsymbol{x}_{k}^{\left(j\right)}\right\}_{k=1}^{\tilde{N}_{j}}\subseteq S^{\left(j\right)}$. We consider the problem:
(3.4) $\mathcal{F}^{\left(j\right)}u\biggr{|}_{S^{\left(j\right)}}$ $=f^{\left(j\right)}\text{, for each $j\in\left\{1,2,\ldots,N_{S}\right\}$},$
where each $\mathcal{F}^{\left(j\right)}$ is a linear differential operator of order at most $p$ and $f^{\left(j\right)}$ is defined in a neighbourhood of each point in $S^{\left(j\right)}$ for each $j$. This form covers all strong-form linear PDEs on manifolds that we may consider. For example, to solve a Laplace-Beltrami problem on a surface $S$, we may set $S^{\left(1\right)}=S^{\left(2\right)}=S$ and $S^{\left(3\right)}=\partial S$. We could then choose $\mathcal{F}^{\left(1\right)}u=\Delta u-\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}u\right)\hat{\boldsymbol{n}}_{S}$, $\mathcal{F}^{\left(2\right)}u=\nabla u-\hat{\boldsymbol{n}}_{S}\left(\nabla u\cdot\hat{\boldsymbol{n}}_{S}\right)$, $f^{\left(1\right)}=f$, and $f^{\left(2\right)}=0$ so that $\Delta_{S}u\biggr{|}_{S}=f$ by Corollary 3. $\mathcal{F}^{\left(3\right)}$ would determine the boundary conditions.
We then have the discretization:
@@
subject to: $\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}b_{n}\left(\mathcal{F}^{\left(j\right)}\phi_{n}\right)\left(\boldsymbol{x}_{k}^{\left(j\right)}\right)=f^{\left(j\right)}\left(\boldsymbol{x}_{k}^{\left(j\right)}\right)\text{, for each $j,k$}.$
@@ 3.5. Analysis for Hermite-Birkhoff Problem
Let $h_{\text{max}}$ be the largest fill distance of each point cloud $S_{\tilde{N}_{j}}^{\left(j\right)}$ in their respective domains $S^{\left(j\right)}$. We can then prove uniform convergence of $\mathcal{F}^{\left(j\right)}\tilde{u}\to f^{\left(j\right)}$; convergence is high-order or super-algebraic with respect to $h_{\text{max}}$ as long as certain smoothness requirements are met. This, or similar arguments with different norms, is enough to prove high-order convergence of $\tilde{u}$ to a solution to the PDE if the PDE is well-posed with respect to the norms in consideration; this is considered in Proposition 11. Proposition 10 will later show convergence to a solution of the PDE directly under weaker conditions, albeit without rate estimates.
Note for the following results that we typically omit the notation for the restriction of a function when the restricted domain is otherwise clear, such as in a norm. For example, we may write $\left\|\tilde{u}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}$ to mean $\left\|\tilde{u}|_{S^{\left(j\right)}}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}$. Noting that if $S^{\left(j\right)}$ is a point, then $\mathcal{F}^{\left(j\right)}\tilde{u}=f^{\left(j\right)}$ on that point as long as $S^{\left(j\right)}_{\tilde{N}_{j}}$ is non-empty, we now cover the case that $S^{\left(j\right)}$ has dimension at least one.
@@
Let $\tilde{b}$ be the solution to (3.5) and let $\tilde{u}=\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{n}\phi_{n}$. Assume $S^{\left(j\right)}$ is of dimension $m_{j}\geq 1$ and can be parametrized by a finite atlas $\left\{\sigma^{\left(j\right)}_{k}\right\}_{k=1}^{M}$ where $\sigma^{\left(j\right)}_{k}:U_{k}^{\left(j\right)}\to S^{\left(j\right)}_{k}$ and each $U_{k}^{\left(j\right)}\subset\mathbb{R}^{m_{j}}$ satisfies an interior cone condition. Let $q\in\mathbb{N}$ and assume that each $\sigma_{k}^{\left(j\right)}$ is $C^{q}$ with bounded derivatives up to order $q$. Also, assume that each $\sigma_{k}^{\left(j\right)}$ has an inverse metric tensor with a bounded norm on $\sigma_{k}^{\left(j\right)}\left(U_{k}^{\left(j\right)}\right)$. Let $j\in\left\{1,2,\ldots,N_{S}\right\}$ be fixed. Assume $f^{\left(j\right)}$ has bounded derivatives up to order $q$ and $F^{(j,q)}d^{-\frac{1}{2}}\in\ell^{2}$, where $F^{(j,q)}:=\left\{\max_{\left|\beta\right|\leq q}\left\|\partial^{\beta}\mathcal{F}^{\left(j\right)}\phi_{n}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right\}_{n=1}^{\infty}$. Then, for small enough $h_{\text{max}}$, there exist constants $\tilde{B}_{j,k,q,\left|\alpha\right|}>0$ such that
$\left\|\partial^{\alpha}\left(\left(\mathcal{F}^{\left(j\right)}\tilde{u}-f^{\left(j\right)}\right)\circ\sigma_{k}^{\left(j\right)}\right)\right\|_{L^{\infty}\left(U_{k}^{\left(j\right)}\right)}\leq\tilde{B}_{j,k,q,\left|\alpha\right|}h_{\text{max}}^{q-\frac{m_{j}}{2}-\left|\alpha\right|}\left\|u\right\|_{d},$
where $u=\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}{b}_{n}\phi_{n}\in\mathcal{H}\left(d\right)$ is a solution to (3.4), $\tilde{C}_{j,q}:=\left\|F^{\left(j,q\right)}d^{-\frac{1}{2}}\right\|_{\ell^{2}}$, and $\alpha$ is a multi-index with $\left|\alpha\right|\in\left\{0,1,2,\ldots,p\right\}$.
@@
Since the derivatives of $\sigma^{\left(j\right)}_{k}$ are bounded, repeated application of chain rule shows there exist constants $A_{k,q}$ so that for each multi-index $\gamma$ of order less than or equal to $q$,
@@
$\ \left\|\partial^{\gamma}\left(\mathcal{F}^{\left(j\right)}\tilde{u}\circ\sigma^{\left(j\right)}_{k}-f^{\left(j\right)}\circ\sigma^{\left(j\right)}_{k}\right)\right\|_{L^{\infty}\left(U^{(j)}_{k}\right)}$
$\leq$ $\ A_{k,q}\max_{\left|\beta\right|\leq q}\left\|\partial^{\beta}\left(\mathcal{F}^{\left(j\right)}\tilde{u}-f^{\left(j\right)}\right)\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}$
$=$ $\ A_{k,q}\max_{\left|\beta\right|\leq q}\left\|\partial^{\beta}\left(\mathcal{F}^{\left(j\right)}\left(\tilde{u}-u\right)\right)\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}$